# Feedback-guided Domain Synthesis with Multi-Source Conditional Diffusion Models for Domain Generalization

### v1

@@ Appendix C Detailed Results
Method Source Domains Caltech101 LabelMe SUN09 VOC2007 Avg. GroupDRO 97.3 $\pm$0.3 63.4 $\pm$0.9 69.5 $\pm$0.8 76.7 $\pm$0.7 76.73 $\pm$0.7 RSC 97.9 $\pm$0.1 62.5 $\pm$0.7 72.3 $\pm$1.2 75.6 $\pm$0.8 77.08 $\pm$0.7 MLDG 97.4 $\pm$0.2 65.2 $\pm$0.7 71.0 $\pm$1.4 75.3 $\pm$1.0 77.23 $\pm$0.8 MTL 97.8 $\pm$0.4 64.3 $\pm$0.3 71.5 $\pm$0.7 75.3 $\pm$1.7 77.23 $\pm$0.7 ERM 98.0 $\pm$0.3 64.7 $\pm$1.2 71.4 $\pm$1.2 75.2 $\pm$1.6 77.33 $\pm$1.8 I-Mixup 98.3 $\pm$0.6 64.8 $\pm$1.0 72.1 $\pm$0.5 74.3 $\pm$0.8 77.38 $\pm$0.7 ERM 97.7 $\pm$0.4 64.3 $\pm$0.9 73.4 $\pm$0.5 74.6 $\pm$1.3 77.50 $\pm$0.8 MMD 97.7 $\pm$0.1 64.0 $\pm$1.1 72.8 $\pm$0.2 75.3 $\pm$3.3 77.45 $\pm$1.1 CDANN 97.1 $\pm$0.3 65.1 $\pm$1.2 70.7 $\pm$0.8 77.1 $\pm$1.5 77.50 $\pm$0.9 ARM 98.7 $\pm$0.2 63.6 $\pm$0.7 71.3 $\pm$1.2 76.7 $\pm$0.6 77.58 $\pm$0.6 SagNet 97.9 $\pm$0.4 64.5 $\pm$0.5 71.4 $\pm$1.3 77.5 $\pm$0.5 77.83 $\pm$0.6 Mixstyle 98.6 $\pm$0.3 64.5 $\pm$1.1 72.6 $\pm$0.5 75.7 $\pm$1.7 77.85 $\pm$0.9 VREx 98.4 $\pm$0.3 64.4 $\pm$1.4 74.1 $\pm$0.4 76.2 $\pm$1.3 78.28 $\pm$0.8 IRM 98.6 $\pm$0.1 64.9 $\pm$0.9 73.4 $\pm$0.6 77.3 $\pm$0.9 78.55 $\pm$0.6 DANN 99.0 $\pm$0.3 65.1 $\pm$1.4 73.1 $\pm$0.3 77.2 $\pm$0.6 78.60 $\pm$0.6 CORAL 98.3 $\pm$0.1 66.1 $\pm$1.2 73.4 $\pm$0.3 77.5 $\pm$1.2 78.83 $\pm$0.7 SWAD 98.8 $\pm$0.1 63.3 $\pm$0.3 75.3 $\pm$0.5 79.2 $\pm$0.6 79.15 $\pm$0.3 ERM 96.91 $\pm$1.4 64.13 $\pm$1.4 71.11 $\pm$1.5 72.78 $\pm$0.9 76.23 $\pm$1.3 ERM + FDS (ours) 99.35 $\pm$0.2 63.71 $\pm$1.7 75.16 $\pm$3.1 79.60 $\pm$1.5 79.46 $\pm$1.6 SWAD 98.65 $\pm$0.2 63.89 $\pm$0.3 74.32 $\pm$1.1 78.61 $\pm$0.6 78.87 $\pm$0.5 SWAD + FDS (ours) 99.47 $\pm$0.2 62.89 $\pm$0.2 76.94 $\pm$0. 79.60 $\pm$1.3 79.73 $\pm$0.5
Table 13: Deatailed Comparison of FDS with the state-of-the-art on OfficeHome dataset. numbers marked are reproduced in-house.
@@
Method Source Domains Art Clipart Product RealWorld Avg. Mixstyle 51.1 $\pm$0.3 53.2 $\pm$0.4 68.2 $\pm$0.7 69.2 $\pm$0.6 60.43 $\pm$0.5 IRM 58.9 $\pm$2.3 52.2 $\pm$1.6 72.1 $\pm$2.9 74.0 $\pm$2.5 64.30 $\pm$2.3 ARM 58.9 $\pm$0.8 51.0 $\pm$0.5 74.1 $\pm$0.1 75.2 $\pm$0.3 64.80 $\pm$0.4 RSC 60.7 $\pm$1.4 51.4 $\pm$0.3 74.8 $\pm$1.1 75.1 $\pm$1.3 65.50 $\pm$1.1 CDANN 61.5 $\pm$1.4 50.4 $\pm$2.4 74.4 $\pm$0.9 76.6 $\pm$0.8 65.73 $\pm$1.3 DANN 59.9 $\pm$1.3 53.0 $\pm$0.3 73.6 $\pm$0.7 76.9 $\pm$0.5 65.85 $\pm$0.7 GroupDRO 60.4 $\pm$0.7 52.7 $\pm$1.0 75.0 $\pm$0.7 76.0 $\pm$0.7 66.03 $\pm$0.8 MMD 60.4 $\pm$0.2 53.3 $\pm$0.3 74.3 $\pm$0.1 77.4 $\pm$0.6 66.35 $\pm$0.3 MTL 61.5 $\pm$0.7 52.4 $\pm$0.6 74.9 $\pm$0.4 76.8 $\pm$0.4 66.40 $\pm$0.5 VREx 60.7 $\pm$0.9 53.0 $\pm$0.9 75.3 $\pm$0.1 76.6 $\pm$0.5 66.40 $\pm$0.6 MLDG 61.5 $\pm$0.9 53.2 $\pm$0.6 75.0 $\pm$1.2 77.5 $\pm$0.4 66.80 $\pm$0.7 ERM 63.1 $\pm$0.3 51.9 $\pm$0.4 77.2 $\pm$0.5 78.1 $\pm$0.2 67.58 $\pm$0.3 I-Mixup 62.4 $\pm$0.8 54.8 $\pm$0.6 76.9 $\pm$0.3 78.3 $\pm$0.2 68.10 $\pm$0.5 SagNet 63.4 $\pm$0.2 54.8 $\pm$0.4 75.8 $\pm$0.4 78.3 $\pm$0.3 68.08 $\pm$0.3 CORAL 65.3 $\pm$0.4 54.4 $\pm$0.5 76.5 $\pm$0.1 78.4 $\pm$0.5 68.65 $\pm$0.4 SWAD 66.1 $\pm$0.4 57.7 $\pm$0.4 78.4 $\pm$0.1 80.2 $\pm$0.2 70.60 $\pm$0.3 ERM 59.54 $\pm$2.1 51.25 $\pm$1.3 73.81 $\pm$0.8 73.78 $\pm$0.2 64.60 $\pm$1.1 ERM + FDS (ours) 60.78 $\pm$1.4 55.90 $\pm$1.6 79.54 $\pm$0.5 80.45 $\pm$0.8 69.17 $\pm$1.1 SWAD 65.86 $\pm$0.9 56.81 $\pm$0.4 78.66 $\pm$0.3 80.04 $\pm$0.2 70.34 $\pm$0.4 SWAD + FDS (ours) 67.34 $\pm$0.8 60.52 $\pm$0.5 82.56 $\pm$0.1 83.62 $\pm$0.3 73.51 $\pm$0.4

### v2

@@ Appendix D Detailed Results
Method Aug. Target Domains Caltech101 LabelMe SUN09 VOC2007 Avg. Standard Methods ERM (baseline) [44] 97.7 $\pm$0.4 64.3 $\pm$0.9 73.4 $\pm$0.5 74.6 $\pm$1.3 77.5 $\pm$0.4 ERM (reproduced) 96.9 $\pm$1.4 64.1 $\pm$1.4 71.1 $\pm$1.5 72.8 $\pm$0.9 76.2 $\pm$1.1 IRM [60] 98.6 $\pm$0.1 64.9 $\pm$0.9 73.4 $\pm$0.6 77.3 $\pm$0.9 78.5 $\pm$0.5 GroupDRO [47] 97.3 $\pm$0.3 63.4 $\pm$0.9 69.5 $\pm$0.8 76.7 $\pm$0.7 76.7 $\pm$0.6 Mixup [61] 98.3 $\pm$0.6 64.8 $\pm$1.0 72.1 $\pm$0.5 74.3 $\pm$0.8 77.4 $\pm$0.6 CORAL [62] 98.3 $\pm$0.1 66.1 $\pm$1.2 73.4 $\pm$0.3 77.5 $\pm$1.2 78.8 $\pm$0.6 MMD [63] 97.7 $\pm$0.1 64.0 $\pm$1.1 72.8 $\pm$0.2 75.3 $\pm$3.3 77.5 $\pm$0.9 DANN [64] 99.0 $\pm$0.3 65.1 $\pm$1.4 73.1 $\pm$0.3 77.2 $\pm$0.6 78.6 $\pm$0.4 MLDG [15] 97.4 $\pm$0.2 65.2 $\pm$0.7 71.0 $\pm$1.4 75.3 $\pm$1.0 77.2 $\pm$0.8 VREx [74] 98.4 $\pm$0.3 64.4 $\pm$1.4 74.1 $\pm$0.4 76.2 $\pm$1.3 78.3 $\pm$0.8 ARM [75] 98.7 $\pm$0.2 63.6 $\pm$0.7 71.3 $\pm$1.2 76.7 $\pm$0.6 77.6 $\pm$0.6 SagNet [65] 97.9 $\pm$0.4 64.5 $\pm$0.5 71.4 $\pm$1.3 77.5 $\pm$0.5 77.8 $\pm$0.5 RSC [23] 97.9 $\pm$0.1 62.5 $\pm$0.7 72.3 $\pm$1.2 75.6 $\pm$0.8 77.1 $\pm$0.5 Mixstyle [20] 98.6 $\pm$0.3 64.5 $\pm$1.1 72.6 $\pm$0.5 75.7 $\pm$1.7 77.9 $\pm$0.5 mDSDI [66] 97.6 $\pm$0.1 66.4 $\pm$0.4 74.0 $\pm$0.6 77.8 $\pm$0.7 79.0 $\pm$0.3 SelfReg [41] 96.7 $\pm$0.4 65.2 $\pm$1.2 73.1 $\pm$1.3 76.2 $\pm$0.7 77.8 $\pm$0.9 Fishr [73] 98.9 $\pm$0.3 64.0 $\pm$0.5 71.5 $\pm$0.2 76.8 $\pm$0.7 77.8 $\pm$0.5 DCAug [59] 98.3 $\pm$0.1 64.2 $\pm$0.4 74.4 $\pm$0.6 77.5 $\pm$0.3 78.6 $\pm$0.5 CDGA [55] 96.3 $\pm$0.7 75.7 $\pm$1.0 72.8 $\pm$1.3 73.7 $\pm$1.3 79.6 $\pm$0.9 ERM + FDS (ours) 98.8 $\pm$0.3 65.6 $\pm$0.9 75.5 $\pm$0.9 79.3 $\pm$1.8 79.8 $\pm$0.5 WA Methods SWAD (baseline) [24] 98.8 $\pm$0.1 63.3 $\pm$0.3 75.3 $\pm$0.5 79.2 $\pm$0.6 79.1 $\pm$0.1 SWAD (reproduced) 98.7 $\pm$0.2 63.9 $\pm$0.3 74.3 $\pm$1.1 78.6 $\pm$0.6 78.9 $\pm$0.5 SelfReg SWA [41] 97.4 $\pm$0.4 63.5 $\pm$0.3 72.6 $\pm$0.1 76.7 $\pm$0.7 77.5 $\pm$0.0 DNA [67] 98.8 $\pm$0.1 63.6 $\pm$0.2 74.1 $\pm$0.1 79.5 $\pm$0.4 79.0 $\pm$0.5 DiWA [68] 98.4 $\pm$0.1 63.4 $\pm$0.1 75.5 $\pm$0.3 78.9 $\pm$0.6 79.1 $\pm$0.2 TeachDCAug [59] 98.5 $\pm$0.1 63.7 $\pm$0.3 75.6 $\pm$0.5 77.0 $\pm$0.7 78.7 $\pm$0.5 SWAD + FDS (ours) 99.5 $\pm$0.2 62.9 $\pm$0.2 76.9 $\pm$0.4 79.6 $\pm$1.3 79.7 $\pm$0.5
Table 14: Leave-one-out accuracy (%) results on the OfficeHome dataset. "Aug." indicates whether advanced augmentation or domain mixing techniques are used. The best results and second-best results are highlighted.
@@
Method Aug. Target Domains Art Clipart Product Real World Avg. Standard Methods ERM (baseline) [44] 61.3 $\pm$0.7 52.4 $\pm$0.3 75.8 $\pm$0.1 76.6 $\pm$0.3 66.5 $\pm$0.3 ERM (reproduced) 59.5 $\pm$2.1 51.3 $\pm$1.3 73.8 $\pm$0.8 73.8 $\pm$0.2 64.6 $\pm$1.1 IRM [60] 58.9 $\pm$2.3 52.2 $\pm$1.6 72.1 $\pm$2.9 74.0 $\pm$2.5 64.3 $\pm$2.2 GroupDRO [47] 60.4 $\pm$0.7 52.7 $\pm$1.0 75.0 $\pm$0.7 76.0 $\pm$0.7 66.0 $\pm$0.7 Mixup [61] 62.4 $\pm$0.8 54.8 $\pm$0.6 76.9 $\pm$0.3 78.3 $\pm$0.2 68.1 $\pm$0.3 CORAL [62] 65.3 $\pm$0.4 54.4 $\pm$0.5 76.5 $\pm$0.1 78.4 $\pm$0.5 68.7 $\pm$0.3 MMD [63] 60.4 $\pm$0.2 53.3 $\pm$0.3 74.3 $\pm$0.1 77.4 $\pm$0.6 66.3 $\pm$0.1 DANN [64] 59.9 $\pm$1.3 53.0 $\pm$0.3 73.6 $\pm$0.7 76.9 $\pm$0.5 65.9 $\pm$0.6 MLDG [15] 61.5 $\pm$0.9 53.2 $\pm$0.6 75.0 $\pm$1.2 77.5 $\pm$0.4 66.8 $\pm$0.7 VREx [74] 60.7 $\pm$0.9 53.0 $\pm$0.9 75.3 $\pm$0.1 76.6 $\pm$0.5 66.4 $\pm$0.6 ARM [75] 58.9 $\pm$0.8 51.0 $\pm$0.5 74.1 $\pm$0.1 75.2 $\pm$0.3 64.8 $\pm$0.4 SagNet [65] 63.4 $\pm$0.2 54.8 $\pm$0.4 75.8 $\pm$0.4 78.3 $\pm$0.3 68.1 $\pm$0.1 RSC [23] 60.7 $\pm$1.4 51.4 $\pm$0.3 74.8 $\pm$1.1 75.1 $\pm$1.3 65.5 $\pm$0.9 Mixstyle [20] 51.1 $\pm$0.3 53.2 $\pm$0.4 68.2 $\pm$0.7 69.2 $\pm$0.6 60.4 $\pm$0.3 mDSDI [66] 68.1 $\pm$0.3 52.1 $\pm$0.4 76.0 $\pm$0.2 80.4 $\pm$0.2 69.2 $\pm$0.4 SelfReg [41] 63.6 $\pm$1.4 53.1 $\pm$1.0 76.9 $\pm$0.4 78.1 $\pm$0.4 67.9 $\pm$0.7 Fishr [73] 62.4 $\pm$0.5 54.4 $\pm$0.4 76.2 $\pm$0.5 78.3 $\pm$0.1 67.8 $\pm$0.5 DCAug [59] 61.8 $\pm$0.6 55.4 $\pm$0.6 77.1 $\pm$0.3 78.9 $\pm$0.3 68.3 $\pm$0.4 CDGA [55] 60.1 $\pm$1.4 54.2 $\pm$0.5 78.2 $\pm$0.6 80.4 $\pm$0.1 68.2 $\pm$0.6 ERM + FDS (ours) 64.6 $\pm$0.2 57.7 $\pm$0.1 80.2 $\pm$0.5 82.0 $\pm$0.4 71.1 $\pm$0.1 WA Methods SWAD (baseline) [24] 66.1 $\pm$0.4 57.7 $\pm$0.4 78.4 $\pm$0.1 80.2 $\pm$0.2 70.6 $\pm$0.2 SWAD (reproduced) 65.9 $\pm$0.9 56.8 $\pm$0.4 78.8 $\pm$0.3 80.0 $\pm$0.2 70.3 $\pm$0.4 SelfReg SWA [41] 64.9 $\pm$0.8 55.4 $\pm$0.6 78.4 $\pm$0.2 78.8 $\pm$0.1 69.4 $\pm$0.2 DNA [67] 67.7 $\pm$0.2 57.7 $\pm$0.3 78.9 $\pm$0.2 80.5 $\pm$0.2 71.2 $\pm$0.1 DiWA [68] 67.3 $\pm$0.2 57.9 $\pm$0.2 79.0 $\pm$0.2 79.9 $\pm$0.1 71.0 $\pm$0.1 TeachDCAug [59] 66.2 $\pm$0.2 57.0 $\pm$0.3 78.3 $\pm$0.1 80.1 $\pm$0.0 70.4 $\pm$0.2 SWAD + FDS (ours) 67.3 $\pm$0.8 60.5 $\pm$0.5 82.6 $\pm$0.1 83.6 $\pm$0.3 73.5 $\pm$0.4
Table 15: Oracle (test-domain validation set) accuracy (%) results on the PACS dataset. "Aug." indicates whether advanced augmentation or domain mixing techniques are used. The best results and second-best results are highlighted.
Method Aug. Target Domains Art Cartoon Photo Sketch Avg. Standard Methods ERM (baseline) [44] 86.5 $\pm$1.0 81.3 $\pm$0.6 96.2 $\pm$0.3 82.7 $\pm$1.1 86.7 $\pm$0.8 ERM (reproduced) 88.6 $\pm$0.9 80.9 $\pm$1.9 98.4 $\pm$0.4 78.4 $\pm$1.2 86.6 $\pm$1.0 IRM 84.2 $\pm$0.9 79.7 $\pm$1.5 95.9 $\pm$0.4 78.3 $\pm$2.1 84.5 $\pm$1.2 GroupDRO 87.5 $\pm$0.5 82.9 $\pm$0.6 97.1 $\pm$0.3 81.1 $\pm$1.2 87.2 $\pm$0.7 Mixup 87.5 $\pm$0.4 81.6 $\pm$0.7 97.4 $\pm$0.2 80.8 $\pm$0.9 86.8 $\pm$0.6 CORAL 86.6 $\pm$0.8 81.8 $\pm$0.9 97.1 $\pm$0.5 82.7 $\pm$0.6 87.1 $\pm$0.7 MMD 88.1 $\pm$0.8 82.6 $\pm$0.7 97.1 $\pm$0.5 81.2 $\pm$1.2 87.3 $\pm$0.8 DANN 87.0 $\pm$0.4 80.3 $\pm$0.6 96.8 $\pm$0.3 76.9 $\pm$1.1 85.3 $\pm$0.6 SagNet 87.4 $\pm$0.5 81.2 $\pm$1.2 96.3 $\pm$0.8 80.7 $\pm$1.1 86.4 $\pm$0.9 RSC 86.0 $\pm$0.7 81.8 $\pm$0.9 96.8 $\pm$0.7 80.4 $\pm$0.5 86.3 $\pm$0.7 Fishr 87.9 $\pm$0.6 80.8 $\pm$0.5 97.9 $\pm$0.4 81.1 $\pm$0.8 86.9 $\pm$0.6 SelfReg 87.9 $\pm$0.5 80.6 $\pm$1.1 97.1 $\pm$0.4 81.1 $\pm$1.3 86.7 $\pm$0.8 CDGA 89.6 $\pm$0.8 85.3 $\pm$0.7 97.3 $\pm$0.3 86.2 $\pm$0.5 89.6 $\pm$0.6 ERM + FDS (ours) 91.1 $\pm$0.3 84.9 $\pm$0.7 97.3 $\pm$0.5 85.6 $\pm$2.3 89.7 $\pm$0.8
Table 16: Oracle (test-domain validation set) accuracy (%) results on the VLCS dataset. "Aug." indicates whether advanced augmentation or domain mixing techniques are used. The best results and second-best results are highlighted.
Method Aug. Target Domains Caltech101 LabelMe SUN09 VOC2007 Avg. Standard Methods ERM (baseline) [44] 97.6 $\pm$0.3 67.9 $\pm$0.7 70.9 $\pm$0.2 74.0 $\pm$0.6 77.6 $\pm$0.5 ERM (reproduced) 98.6 $\pm$0.2 68.6 $\pm$0.7 73.6 $\pm$1.7 78.6 $\pm$1.2 79.8 $\pm$0.4 IRM 97.3 $\pm$0.2 66.7 $\pm$0.1 71.0 $\pm$2.3 72.8 $\pm$0.4 77.0 $\pm$0.8 GroupDRO 97.7 $\pm$0.2 65.9 $\pm$0.2 72.8 $\pm$0.8 73.4 $\pm$1.3 77.5 $\pm$0.6 Mixup 97.8 $\pm$0.4 67.2 $\pm$0.4 71.5 $\pm$0.2 75.7 $\pm$0.6 78.1 $\pm$0.4 CORAL 97.3 $\pm$0.2 67.5 $\pm$0.6 71.6 $\pm$0.6 74.5 $\pm$0.0 77.7 $\pm$0.4 MMD 98.8 $\pm$0.0 66.4 $\pm$0.4 70.8 $\pm$0.5 75.6 $\pm$0.4 77.9 $\pm$0.3 DANN 9.0 $\pm$0.2 66.3 $\pm$1.2 73.4 $\pm$1.4 80.1 $\pm$0.5 79.7 $\pm$0.8 SagNet 97.4 $\pm$0.3 66.4 $\pm$0.4 71.6 $\pm$0.1 75.0 $\pm$0.8 77.6 $\pm$0.4 RSC 98.0 $\pm$0.4 67.2 $\pm$0.3 70.3 $\pm$1.3 75.6 $\pm$0.4 77.8 $\pm$0.6 Fishr 97.6 $\pm$0.7 67.3 $\pm$0.5 72.2 $\pm$0.9 75.7 $\pm$0.3 78.2 $\pm$0.6 SelfReg 98.2 $\pm$0.3 63.9 $\pm$0.8 72.2 $\pm$0.1 75.5 $\pm$0.4 77.5 $\pm$0.2 CDGA 96.6 $\pm$0.7 75.5 $\pm$1.9 73.6 $\pm$1.1 77.8 $\pm$1.0 80.9 $\pm$1.2 ERM + FDS (ours) 99.5 $\pm$0.1 68.7 $\pm$0.3 77.4 $\pm$0.7 82.6 $\pm$0.1 82.0 $\pm$0.1
Table 17: Oracle (test-domain validation set) accuracy (%) results on the OfficeHome dataset. "Aug." indicates whether advanced augmentation or domain mixing techniques are used. The best results and second-best results are highlighted.
Method Aug. Target Domains Art Clipart Product Real World Avg. Standard Methods ERM (baseline) [44] 61.7 $\pm$0.7 53.4 $\pm$0.3 74.1 $\pm$0.4 76.2 $\pm$0.6 66.4 $\pm$0.5 ERM (reproduced) 64.0 $\pm$0.9 53.7 $\pm$1.1 77.1 $\pm$0.3 78.8 $\pm$0.4 68.4 $\pm$0.3 IRM 56.4 $\pm$3.2 51.2 $\pm$2.3 71.7 $\pm$2.7 72.7 $\pm$2.7 63.0 $\pm$2.7 GroupDRO 60.5 $\pm$1.6 53.1 $\pm$0.3 75.5 $\pm$0.3 75.9 $\pm$0.7 66.3 $\pm$0.7 Mixup 63.5 $\pm$0.2 54.6 $\pm$0.4 76.0 $\pm$0.3 78.0 $\pm$0.7 68.0 $\pm$0.4 CORAL 64.8 $\pm$0.8 54.1 $\pm$0.9 76.5 $\pm$0.4 78.2 $\pm$0.4 68.4 $\pm$0.6 MMD 60.4 $\pm$1.0 53.4 $\pm$0.5 74.9 $\pm$0.1 76.1 $\pm$0.7 66.2 $\pm$0.6 DANN 60.6 $\pm$1.4 51.8 $\pm$0.7 73.4 $\pm$0.5 75.5 $\pm$0.9 65.3 $\pm$0.9 SagNet 62.7 $\pm$0.5 53.6 $\pm$0.5 76.0 $\pm$0.3 77.8 $\pm$0.1 67.5 $\pm$0.4 RSC 61.7 $\pm$0.8 53.0 $\pm$0.9 74.8 $\pm$0.8 76.3 $\pm$0.5 66.5 $\pm$0.8 Fishr 63.4 $\pm$0.8 54.2 $\pm$0.3 76.4 $\pm$0.3 78.5 $\pm$0.2 68.1 $\pm$0.4 SelfReg 64.2 $\pm$0.6 53.6 $\pm$0.7 76.7 $\pm$0.3 77.9 $\pm$0.5 68.1 $\pm$0.3 CDGA 61.1 $\pm$1.1 55.9 $\pm$1.0 78.2 $\pm$0.8 79.8 $\pm$0.2 68.8 $\pm$0.8 ERM + FDS (ours) 65.3 $\pm$0.8 58.4 $\pm$0.8 81.2 $\pm$0.2 82.4 $\pm$0.6 71.8 $\pm$0.9